OpenStudy (anonymous):

Which set of points would result in a slope of 1/4? (8, 2) and (4, 1) (8, 2) and (1, 4) (2, 8) and (4, 1) (2, 8) and (1, 4)

jimthompson5910 (jim_thompson5910):

Hint: use the slope formula \[\Large m = \frac{y_{2}-y_{1}}{x_{2}-x_{1}}\]

jimthompson5910 (jim_thompson5910):

For example, the slope through (8,2) and (1,4) is... \[\Large m = \frac{y_{2}-y_{1}}{x_{2}-x_{1}}\] \[\Large m = \frac{4-2}{1-8}\] \[\Large m = \frac{2}{-7}\] \[\Large m = -\frac{2}{7}\] So choice B isn't the answer